# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

### v1

@@ Abstract
Semi-supervised learning is designed to help reduce the cost of the manual labelling process by exploiting the utilisation of useful features from a large quantity of unlabelled data during training. Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this. However, most of the existing semi-supervised learning methods still lack efficient perturbation methods to promote the precision and diversity of features (or pseudo labels) during training. In order to fill this gap, we propose DiverseNet architectures which explore multi-head and multi-model semi-supervised learning algorithms by simultaneously promoting precision and diversity during training. The two proposed methods of DiverseNet, namely the DiverseHead and DiverseModel, significantly achieve the highest semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods. Meanwhile, the proposed DiverseHead architecture is relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all utilised data sets.
Index Terms:
Semi-supervised Learning, Consistency Learning, Land Cover Classification
I Introduction
@@
Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade. Specifically, with its great success in semantic segmentation, automated land cover classification approaches have been significantly improved because of the high similarity between the two tasks [vali2020deep]. Nevertheless, supervised learning necessitates a substantial, meticulously labelled dataset. In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations proves to be a time-consuming, labour-intensive, and costly endeavour. While the field of computer vision boasts numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge. This is primarily because of the significant disparities between natural images and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, which often possesses unconventional image geometries, making the process non-intuitive. Fortunately, thanks to open-access remote sensing sources, a large amount of unlabelled imagery can be accessed easily and freely. Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets [wang2022semi].
@@
In computer vision, semi-supervised learning has become a trending research direction to eliminate the labour-intensive and expensive annotation stages mentioned above. It leads to competitive performance compared to supervised learning in applications such as image classification [zhang2021flexmatch, huang2021embedding, wang2021semi2] and segmentation [hu2021semi, wang2022semi_u2pl, wang2021semi], which just use a part of labelled data and a large amount of unlabelled data. Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step. Since pseudo labels are involved in the model training stage, the accuracy of pseudo labels is regarded as a key factor affecting the performance of semi-supervised learning approaches. However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough. In this case, augmenting the pseudo-labels to increase diversity has emerged as another key research direction within the semi-supervised learning domain. A greater diversity aids in constructing more comprehensive pseudo-labels, which, in turn, indirectly enhances the quality of these labels and the training effectiveness of the model.
@@
Figure 1: Two kinds of Pseudo label generation methods based on multiple heads (a) DiverseHead and multiple models (b) DiverseModel.
In order to empower networks to harness the unexplored potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning. [french2019semi, filipiak2021n]. Specifically, consistency regularization methods rely on the assumption of smoothing, which posits that if two points reside in a high-density region and are close to each other, their corresponding labels should be the same or consistent [chen2010semi]. In practice, these consistency regularization methods achieve this assumption by forcing networks to produce consistent predictions for unlabeled inputs that undergo diverse perturbations.
The strategies for the above-mentioned perturbations can be categorised into three groups, namely input, feature, and network perturbations. Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process. In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs. Widely used input perturbation methods employed in semi-supervised learning include Cutout[devries2017improved], MixUp [zhang2017mixup], and CutMix [yun2019cutmix]. Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them [french2019semi]. However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples. This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels. Cross-consistency training (CCT) [ouali2020semi] presents a feature perturbation method that introduces noise to both low- and high-level features. These perturbed features are then fed into multi-decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders. However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data. Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process. Lastly, the network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS)[chen2021semi], confidence-guided semi-supervised learning (CGSSL)[ma2023confidence], mean teacher (MT), [tarvainen2017mean], CCT [ouali2020semi], guided collaborative training (GCT) [ke2020guided] and cross teaching between CNN and transformer [luo2022semi]. Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) utilized.
I-A Contributions
The preceding conversation underscores that, although perturbation techniques have their advantages, they also bring inherent drawbacks. Therefore, the creation of methods based on perturbations is a complex task, necessitating the exploration of more sophisticated approaches to augment the potential of semi-supervised learning frameworks by incorporating advanced perturbation methods. This paper proposes perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of two novel feature and decision diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery. First, we develop a lightweight semi-supervised semantic segmentation network called DiverseHead based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF). In addition, we also propose a high-performance multi-model structure called DiverseModel for scenarios equipped with high-memory computational resources. Specifically, the contributions of this work are as follows:
@@ I-A Contributions
    We promote utilisation of multiple decision heads to provide a semi-supervised learning structure, called DiverseHead. This structure is inspired by Bagging (also called bootstrap aggregating) and provides diversity for parameters and features. The main source of perturbation is from the dynamic choosing heads to be updated for each iteration. This architecture does not distinguish the main or auxiliary decoder. Also, all decoders are trained by both supervised and unsupervised losses, which comes with the advantages of high-quality perturbation and diversity of semi-supervised segmentation tasks for remote sensing imagery, whilst providing a relatively lightweight architecture.
@@
    For the purposes of creating perturbation in diversifying decisions, we ustilised two important techniques, namely ’dynamic freezing’ and ’dropout’. We make these techniques fit with DiverseHead semi-supervised learning architecture for the purpose of diversifying the parameters of the network. The two perturbation methods differ from input/feature perturbation methods, in which the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
@@
    A previously proposed architecture DiverseModel [ma2023confidence] has been extended for detailed evaluation and in a detailed comparison study for various semantic segmentation data sets. This architecture is based on consistency learning among multiple neural networks with distinct weights.
@@
The rest of the paper is organised as follows: subsection I-B discusses the related work on semi-supervised semantic segmentat ion in remote sensing and some basic knowledge on ensemble machine learning whilst in Section II and III, the proposed algorithms DiverseHead and DiverseModel are presented. Section IV describes the segmentation dataset of remote sensing imagery. The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section V. Section VI concludes the paper with a summary.
I-B Related Work
@@
Semantic segmentation is one of the well-studied and wide-spread research areas in remote sensing society, for applications such as land cover classification/mapping [dong2019land], building change detection [zheng2022hfa], road extraction [ghandorh2022semantic], and marine debris detection [booth2023high]. Also, the number of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas. Fully convolutional networks (FCNs) [long2015fully] is one group of the most widespread networks for segmentation tasks. They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision. Following the FCNs’ success, SegNet [badrinarayanan2017segnet] and UNet [ronneberger2015u] adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder. Alternatively, PSPNet [zhao2017pyramid] introduced a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing. DeepLab architecture [chen2017deeplab] proposed a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernel via setting different dilation rates. DeepLab has then been extended to DeepLabv3+ [chen2018encoder] as an improved hybrid approach that combines an encoder-decoder structure and the ASPP. Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries[xia2021cloud, zhang2019multi]. DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
@@
Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data. Self-training [lee2013pseudo] (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classification [kim2020distribution] and segmentation [zhu2021improving] purposes that involves the generation of pseudo labels through model predictions. These pseudo-labels are subsequently utilized to retrain the model. Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels. Another widely developed semi-supervised learning approach called consistency regularization [french2019semi] is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations. CCT [ouali2020semi] employs an encoder-decoder architecture with multiple auxiliary decoders. These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder. They calculate MSE loss between predictions of the main decoder and each auxiliary decoder without creating pseudo labels. It is worth noting that the unsupervised loss is not used to supervise the main decoder. Following CCT, subsequent methods like GCT [ke2020guided] and CPS [chen2021semi] are proposed to introduce network perturbation for consistency regularization. Specifically, they use two same network structures but with different weight initialization. CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector. Although the network perturbation achieved through different weight initialization in the previous two works shows a good ability to generate diversity, we believe that employing different networks could yield even greater diversity.
@@
Ensemble machine learning is a concept that employs multiple learners and combines their predictions [sewell2008ensemble]. Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor [breiman1996bagging]. Specifically, Bagging creates sample subsets by randomly selecting from the training dataset and subsequently utilizes these acquired subsets to train the foundational models for integration. When forecasting a numerical result, the aggregation involves taking an average of the versions, while predicting a class, it relies on a majority vote. Based on the concept, bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms in terms of classification and regression applications [ren2016ensemble].
II Cross-head Supervision (DiverseHead)

### v2

@@ Abstract
Semi-supervised learning aims to help reduce the cost of the manual labelling process by leveraging valuable features extracted from a substantial pool of unlabeled data alongside a limited set of labelled data during the training phase. Since pixel-level manual labelling in large-scale remote sensing imagery is expensive, semi-supervised learning becomes an appropriate solution to this. However, most of the existing consistency learning frameworks based on network perturbation are very bulky. There is still a lack of lightweight and efficient perturbation methods to promote the diversity of features and the precision of pseudo labels during training. In order to fill this gap, we propose DiverseNet which explores multi-head and multi-model semi-supervised learning algorithms by simultaneously enhancing precision and diversity during training. The two proposed methods in the DiverseNet family, namely DiverseHead and DiverseModel, both achieve the better semantic segmentation performance in four widely utilised remote sensing imagery data sets compared to state-of-the-art semi-supervised learning methods. Meanwhile, the proposed DiverseHead architecture is simple and relatively lightweight in terms of parameter space compared to the state-of-the-art methods whilst reaching high-performance results for all the tested data sets.
keywords:
Semi-supervised Learning, Consistency Learning, Semantic Segmentation
[…]
organization=School of Computer Science and Informatics, Cardiff University,city=Cardiff, postcode=CF24 4HJ, country=United Kingdom
1 Introduction
@@
Supervised deep learning has gradually become the dominant technique in computer vision, especially during the last decade. Specifically, with its great success in semantic segmentation in computer vision, automated land cover classification approaches have been significantly improved  (Vali et al., 2020). Nevertheless, supervised learning necessitates a substantial and meticulously labelled dataset. In the case of extensive remote sensing data, such as satellite imagery and drone-captured images in complex terrains, acquiring pixel-wise expert annotations is a time-consuming, labour-intensive, and costly endeavour. While the field of computer vision provides numerous well-annotated datasets, transferring deep learning models trained on these datasets to the remote sensing domain is a formidable challenge. This is primarily due to the significant differences between images generally used in computer vision and remote sensing data, including hyperspectral and synthetic aperture radar (SAR) imagery, in which images often possess unconventional and non-intuitive characteristics. Fortunately, thanks to open-access remote sensing data sources, a large amount of unlabelled imagery can be accessed easily and freely. Thus, only using a limited amount of labelled data and exploiting the benefits of a large amount of unlabelled data could be a feasible solution for solving the problem of the lack of labelled remote sensing datasets (Wang et al., 2022a; Yang et al., 2023; Wang et al., 2023).
@@
In computer vision, semi-supervised learning has become a popular research direction to eliminate the labour-intensive and expensive annotation stages mentioned above. It leads to competitive performance compared to supervised learning in applications such as image classification (Zhang et al., 2021; Huang et al., 2021; Wang et al., 2021b) and segmentation (Hu et al., 2021; Wang et al., 2022b, 2021a), which exploit just a part of labelled data and a large amount of unlabelled data. Specifically, mainstream semi-supervised learning methods exploit “pseudo” labels that are fake and/or auto-generated label information obtained as a result of an unsupervised step. Since pseudo labels are involved in the model training stage, the accuracy of pseudo labels is regarded as a key factor affecting the performance of semi-supervised learning approaches. However, the accuracy of pseudo labels is not always satisfactory especially when labelled training data is not complete and sufficient enough. Meanwhile, increasing diversity of the pseudo-labels has become another key research direction within the semi-supervised learning domain, especially for consistency regularization (French et al., 2019). This is because a greater diversity of pseudo-labels enhances the effectiveness of training deep learning networks.
In order to empower networks to harness the potentials within unlabeled data, a notable technique known as consistency regularization has emerged as a widely adopted method for semi-supervised learning. (French et al., 2019; Filipiak et al., 2021). Specifically, consistency regularization methods are built up on the theory of assumption of smoothness, which posits that if two points reside in a high-density region of feature space and are close to each other, their corresponding labels should be the same or consistent (Chen and Wang, 2010; Luo et al., 2018; French et al., 2019). In practice, these methods for consistency regularization ensure adherence to this assumption by compelling networks to generate predictions that remain consistent across modified iterations of unlabeled inputs.
The strategies for the aforementioned perturbations can be categorised into three groups, namely input, feature, and network perturbations. Input perturbation is a method that involves making adjustments or alterations to the input data as part of the training process. In this way, the semi-supervised learning approach promotes the consistency of predictions made in response to these modified inputs. Widely used input perturbation methods employed in semi-supervised learning include Cutout (DeVries and Taylor, 2017), MixUp (Zhang et al., 2017), and CutMix (Yun et al., 2019). Among them, CutMix stands out by combining aspects of MixUp and CutOut techniques, demonstrating superior performance compared to them (French et al., 2019). However, input perturbation introduces artificial noise to the data, which might lead to incorrect or noisy labels when generating pseudo-labels for unlabeled examples. This might negatively impact the training process by providing incorrect guidance to the model when using pseudo labels. Cross-consistency training (CCT) (Ouali et al., 2020) presents a feature perturbation method that introduces noise to both low- and high-level features. These perturbed features are then fed into multiple decoders to generate multiple outputs, followed by the enforcement of consistency among the outputs obtained from different decoders. However, feature perturbation alters the intrinsic characteristics and relationships between features, which also might lead to the creation of artificial feature representations that do not accurately reflect the underlying patterns in the original data. Consequently, the perturbed features could result in incorrect or noisy labels when generating pseudo-labels for unlabeled examples, which negatively impacts the learning process. Lastly, network perturbation uses different networks or the same network with different initialization weights to promote diversity of predictions, such as cross pseudo supervision (CPS) (Chen et al., 2021), confidence-guided semi-supervised learning (CGSSL) (Ma et al., 2023), mean teacher (MT), (Tarvainen and Valpola, 2017), CCT (Ouali et al., 2020), guided collaborative training (GCT) (Ke et al., 2020) and cross teaching between CNN and transformer (Luo et al., 2022). Compared to the previous two perturbation methods, although network perturbation methods introduce perturbations created by the model itself instead of artificial noise, they generally require rapidly increased computational resources due to the increased number of networks (whole or their inner stages) that are used.
@@
[…]
1.1 Contributions
The preceding overview underscores that, although previous perturbation techniques in semi-supervised learning have their advantages, they also bring inherent drawbacks which could be a key consideration affecting the performance in semi-supervised learning. In addressing the challenges posed by perturbation-based semi-supervised learning and its complexity, it becomes imperative to explore more sophisticated and efficient approaches. These approaches aim to enhance the capabilities of semi-supervised learning frameworks by integrating advanced perturbation methods. This paper proposes two perturbation-based semi-supervised network architectures, coined as DiverseNet, which consist of decision (DiverseHead) and novel feature (DiverseModel) diversifying network architectures for the purposes of semi-supervised semantic segmentation of remote sensing imagery. The proposed lightweight semi-supervised semantic segmentation network called DiverseHead is shown in Figure 1-(a) based on multiple decision heads to promote precision and diversity of pseudo labels whilst using two novel perturbation methods of dropout (DT) and dynamic freezing (DF). In addition, we also analysed a previously proposed high-performance multi-model structure called DiverseModel (Ma et al., 2023) as shown in Figure 1-(b) for scenarios equipped with high-memory computational resources. Specifically, the contributions of this work are as follows:
@@ 1.1 Contributions
    We proposed a simple but efficient semi-supervised learning architecture, promoting the utilisation of multiple decision heads called DiverseHead. This structure is inspired by bagging (also called bootstrap aggregating) and provides diversity for parameters and features. Since each head only consists of a couple of convolutional layers instead of a whole network, the architecture is relatively lightweight and adaptable to various segmentation networks.
@@
    In order to provide efficient pseudo labels in the training stage, we provide a comprehensive voting mechanism based on the proposed DiverseHead structure. The voting mechanism not only uses the joint output from multiple heads but also individually considers pseudo labels generated from these outputs, which are called Mean Voting and Max Voting in this paper, respectively.
@@
    For the purposes of creating perturbation in diversifying decisions, we used two important techniques, namely ‘dynamic freezing’ and ‘dropout’. We make these techniques fit with the DiverseHead semi-supervised learning architecture to diversify the parameters of the network. The two perturbation methods differ from input/feature perturbation methods since the source of feature diversity is from parameter diversity instead of introducing artificial noise to data or features.
  4. 4.
    Further analysis and more detailed evaluation has been carried out on a previously proposed architecture DiverseModel (Ma et al., 2023) which is based on consistency learning among multiple neural networks with distinct weights. We provide a detailed comparison study for various semantic segmentation datasets in this paper. Also, we use Grad-CAM (Selvaraju et al., 2020) to verify the observation that different networks exhibit varied attention to the same input, culminating in the combined pseudo label derived from the multiple model predictions showcasing superior quality.
@@
The rest of the paper is organised as follows: Section 2 discusses the related work on semi-supervised semantic segmentation in remote sensing and some basic knowledge on ensemble machine learning whilst in Section 3 and 4, the proposed algorithms DiverseHead and DiverseModel are presented. Section 5 describes the utilised segmentation dataset of remote sensing imagery. The experimental setting and the qualitative and quantitative analyses of the experimental results are performed in Section 6. Section 7 concludes the paper with a summary.
2 Related Work
@@
Semantic segmentation is a well-studied and wide-spread research topic in remote sensing, incorporated within applications such as land cover classification/mapping (Dong et al., 2019), building change detection (Zheng et al., 2022), road extraction (Ghandorh et al., 2022), and marine debris detection (Kikaki et al., 2022; Booth et al., 2023). The volume of research on semantic segmentation in remote sensing in the literature is rapidly increasing with the success of deep learning in computer vision, due to the strong task-similarity between two areas. Fully convolutional networks (FCNs) (Long et al., 2015) is one group of the most widespread networks for segmentation tasks. They have made a considerable contribution to various segmentation tasks either in remote sensing or computer vision. Following the FCNs’ success, SegNet (Badrinarayanan et al., 2017) and UNet (Ronneberger et al., 2015) adopt a symmetrical encoder-decoder structure with skip connections, leveraging multi-stage features within the encoder. Alternatively, PSPNet (Zhao et al., 2017) introduces a pyramid pooling structure that helps provide a global contextual understanding for pixel-level scene parsing. DeepLab architecture (Chen et al., 2017) proposes a new type of convolution called atrous convolution and an atrous spatial pyramid pooling (ASPP) operation to enable the network to have the ability to control the spatial perception range of convolution kernels by setting different dilation rates. DeepLab has been extended to DeepLabv3+ (Chen et al., 2018) with an improved hybrid approach that combines an encoder-decoder structure and the ASPP. Within DeepLabv3+, the ASPP module is incorporated following a ResNet backbone, enabling the network to delve into deeper layers and extract high-level features, which is helpful for accurately segmenting pixels around boundaries (Xia et al., 2021; Zhang et al., 2019). DeepLabv3+ is one of the most used networks in the literature for semi-supervised learning segmentation in the computer vision area.
@@
Semi-supervised learning aims to alleviate the need for expensive annotation work and improve the models’ performances by making use of unlabelled data. Self-training (Lee et al., 2013) (also known as pseudo labelling) represents one of the primitive semi-supervised learning strategies for both classifications (Kim et al., 2020) and segmentation (Zhu et al., 2021), and involves the generation of pseudo labels through model predictions. These pseudo-labels are subsequently used to retrain the model. Various variations of self-training have been proposed focusing on determining the appropriate pseudo labels. Another widely developed semi-supervised learning approach called consistency regularization (French et al., 2019) is to force networks to give consistent predictions for unlabeled inputs that undergo diverse perturbations. CCT (Ouali et al., 2020) employs an encoder-decoder architecture with multiple auxiliary decoders. These decoders introduce diversity in the output by feature perturbations specific to each auxiliary decoder. They calculate MSE loss between predictions of the main decoder and each auxiliary decoder without creating pseudo labels. It is worth noting that the unsupervised loss is not used to supervise the main decoder. Following CCT, subsequent methods like GCT (Ke et al., 2020) and CPS (Chen et al., 2021) have been proposed to introduce network perturbation for consistency regularization. Specifically, they use two same network structures but with different weight initialization. CPS differs from GCT by employing pseudo labels to enforce consistency, whereas GCT achieves consistency regularization by utilizing predictions of networks cooperating with a flow detector. Although the network perturbation was achieved through different weight initialization in GCT and CPS, the ability to generate diversity is still limited. ICNet (Wang et al., 2022a) was proposed to use teacher networks for increasing the model difference to enhance the diversity of pseudo labels based on an iterative contrast network. Nonetheless, these methods consistently rely on the same network architecture. To introduce more diversity, we advocate the use of different model architectures. By leveraging various networks, we can obtain distinct and complementary features from these models even with the same input data. Thus, the proposed DiverseModel differs from the aforementioned methods, exploring different networks in parallel to generate more diversity of pseudo labels to improve training effectiveness.
@@
Ensemble machine learning is a concept that employs multiple learners and combines their predictions (Sewell, 2008). Bagging, a form of ensemble learning, is a technique aimed at reducing prediction variance by creating multiple iterations of a predictor and then utilizing them to form an aggregated predictor (Breiman, 1996). Specifically, bagging creates sample subsets by randomly selecting from the training data set and subsequently utilises these acquired subsets to train the foundational models for integration. When forecasting a numerical result, the aggregation involves taking an average of the versions, while for predicting a class it relies on a majority vote. Bagging is a commonly employed approach for enhancing the robustness and precision of machine learning algorithms for classification and regression (Ren et al., 2016).
3 DiverseHead (Cross-head Supervision)